a ball is thrown horizontally with an initial speed of 10ms from an 80 m cliff.how long does it take to reach the ground?

1
Expert's answer
2022-06-17T08:14:50-0400

Given:

"h=80\:\rm m"

"g=9.8\:\rm m/s^2"

"v_0=10\:\rm m/s"


The time of motion is equal to the time of free falling:

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2*80}{9.8}}=4.0\:\rm s"